OpenStudy (anonymous):

An airplane flying at an altitude of 40,000 feet passes directly over a fixed object on the ground. One minute later, the angle of depression of the object is 40°. Approximate the speed of the airplane to the nearest mile per hour.

OpenStudy (anonymous):

tan40=40,000/d d=40000/tan40=47,670.14 feet now speed v =d/t 47670.14 feet 1 mile 60 min v= ------------ x ----------- x ----- 1 min 5280 feet 1 hour v= 541.71 mph approx ans..